# Taming Data and Transformers for Audio Generation

### v1

@@ 4.2 Text-2-Audio Generation
Results. In Tab. 5, we report evaluation results. Our method achieves superior performance compared to the state-of-the-art methods in terms of IS, FAD, and $\text{CLAP}_{LAION}$ scores, marking an improvement of $22.7\%$, $15.7\%$, and $13.5\%$, respectively. This shows that GenAu can produce high audio quality and achieve better semantic alignment with the conditioning text. We also find that, while the smaller model can learn the target distribution better, as noted by a lower FD and FAD score, the larger model achieves better IS and $\text{CLAP}_{LAION}$, indicating that more scaling to the audio generator might result in even more enhancement of the audio quality and its alignment with the input prompt.
Ablations. We evaluate our main method variations in Tab. LABEL:tab:audio_generation_ablation. We first test the effect of scaling up the dataset size. When compared with training the small model on AC only, we observe that training the small model on the full dataset without re-captioning produces improvements in IS but degrades both FD and FAD. Instead, when training on the full dataset after re-captioning, we notice consistent and significant improvements on all metrics for both the small and large models. This demonstrates that scaling data introduces significant improvements, but only if the annotation quality is adequately curated. Second, scaling up the model size gives noticeable improvements, especially in audio quality whether the model was trained on the full dataset or only AC. Finally, we compare our small model against a U-Net [93] baseline at a similar computational complexity. We notice significant improvements across all metrics enabled by our scalable transformer architecture.
@@ 5 Conclusion
We take a holistic approach to improving the quality of existing audio generators. starting by addressing the scarcity of large-scale captioned audio datasets, we build a state-of-the-art audio captioning method, AutoCap, which leverages audio metadata to collect a dataset of $761,000$ annotated audio clips. We then built a latent diffusion model based on a scalable transformer architecture which we train on our re-captioned dataset to obtain AutoCap, a state-of-the-art model for audio generation. Our approach not only enhances audio generation but also broadens potential applications. Since AutoCap is fully automated, it can be used to obtain an audio dataset that order of magnitudes larger than the current available ones. Additionally, since AutoCap could be used for captioning current video datasets, it enables novel applications text-to-audio-video joint generation, a more natural and desired choice of video generation. Moreover, our scalable audio generator, GenAu, could be extended to other domains such as speech and music.
Appendix A Limitations.
Audio captioning. Sounds emitted by different objects may be remarkably similar (i.e. a waterfall and heavy rain, or a can and a motorcycle engine). In these situations, if metadata does not have sufficient information, the audio captioning model may not be able to correctly disambiguate sounds. In addition, we find the model may not capture accurately time dependencies between different sounds in the audio and correctly disentangle foreground from background sounds. Additionally, since our captioning model is only finetuned on AudioCap, and Audicaps has a limited vocabulary (less than 1000 words), our model also tends to produce repetitive words and captions.
@@
Audio generation. As we train the model to generate natural sound effects, our audio generation model underperforms in more specific domains where more targeted models might excel such as in music generation or text-to-speech synthesis. Additionally, even though we train on a large dataset, the vocabulary of the paired texts is limited making the model struggle to generate accurate audio for long and descriptive prompts.
Appendix B Training Details
We train our audio generation model using the LAMB optimizer with a learning rate of $0.005$ using a cosine schedule, a weight decay of $0.1$ and a dropout factor of $0.1$. For the small model we train for 100k steps using a batch size of 512. For the large model variant, we train for 50k steps using a batch size of 960. For the last steps of model training, rather than training on the full dataset, we fintune it on AudioCaps only as it constitutes the dataset with captions of highest quality.
Appendix C Additional Evaluation
In this section, we present additional evaluation details and results which are complemented by our Website.
C.1 Evaluation details
@@
While the established practice in the evaluation of audio captioning methods is to report the results on the test set using the checkpoint that performs best on the validation subset, prior work [136, 49] reported high instability of the metrics on the validation subset and weak correlation between the validation and test performance, making the model’s results vary significantly for different seeds. To alleviate this, ENCLAP [49] selects around five best performing validation checkpoints and report their best results on the test set. CNext-trans [136] uses FENSE score to pick the best validation checkpoint. Our model, thanks to the two-stage training paradigm, significantly reduces this instability and we observe steady performance gains as training progresses. Therefore, we report the results at convergence, specifically after 20 epochs.
C.2 Additional Captioning Evaluation
In Tab. 7 we show qualitative results of the captions produced by our method and compare it with state-of-the-art AAC methods. See the Website for qualitative results accompanied by the original audio. While ENCLAP [49] and CoNeTTE [136] tend to produce shord captions, our method produces the most descriptive captions, capturing the most amount of elements from the ground truth, an important capability to allow high-quality audio generation [98].
@@ C.2 Additional Captioning Evaluation
CoNeTTE [136] A man is speaking and wind is blowing
ENCLAP [49] A man is speaking and wind is blowing
@@
CoNeTTE [136] A man speaking with birds chirping in the background.
ENCLAP [49] Birds are chirping and a man speaks
@@
CoNeTTE [136] A telephone ringing followed by a beep.
ENCLAP [49] A telephone dialing followed by a series of electronic beeps
@@
CoNeTTE [136] A train horn blows and a steam whistle is blowing
ENCLAP [49] A train running on railroad tracks followed by a train horn blowing as wind blows into a microphone
@@
CoNeTTE [136] A woman speaking followed by a door opening and closing.
ENCLAP [49] A young girl speaks followed by a loud bang
@@
CoNeTTE [136] A baby is laughing and people are talking.
ENCLAP [49] A baby laughs and splashes in water
@@
CoNeTTE [136] A dog is barking and a person is walking.
ENCLAP [49] Birds chirp and a dog barks
@@
CoNeTTE [136] A toilet is flushed and water is running.
ENCLAP [49] A faucet is turned on and runs
C.3 Additional Audio Generation Evaluation
In Tab. 9, we ablate the 1D-VAE bottleneck size in terms of reconstruction loss and performance of a subsequently trained latent audio diffusion model, in terms of FAD, FD and IS. Similarly to the phenomenon observed in the image and video generation domain [35, 23], we observe that larger number of channels allocated to the latent space result in lower reconstruction losses, but make the latent space more complex, hindering generation quality. We adopt 64 1D-VAE channels for all our experiments.
@@ C.3 Additional Audio Generation Evaluation
Appendix D Computational resources
We train our models on A100 80GB GPUs. Our capioning model is trained 9 hours on 8 A100 80GB GPUs. Our largest audio generation model is trained for 48h on 48 A100 80GB GPUs.

### v2

@@ 4.2 Text-2-Audio Generation
Results. In Tab. 4, we report evaluation results. Our method achieves superior performance compared to the state-of-the-art methods in terms of IS, FAD, $\text{CLAP}_{MS}$ and $\text{CLAP}_{LAION}$ scores, marking an improvement of $22.7\%$, $4.7\%$, $3.6\%$, and $13.5\%$, respectively. This shows that GenAu can produce high audio quality and achieve better semantic alignment with the conditioning text.
Data scaling. We consider two key aspects: data quality and quantity. First, in Tab. 5 ($2^{nd}$ fow), we show that GenAu-L trained with AutoReCap is generally favoured over training only with AudioCaps (AC). This is confirmed in Tab. 3 ($1^{st}$ vs $2^{nd}$ row), where increasing the dataset size significantly boosts the results across all metrics, improving IS by $56.3\%$. Additionally, we show ($3^{rd}$ vs $4^{th}$ row) that using AutoCap to recaption the dataset significantly enhances the results over all metrics, confirming the importance of data quality. Interestingly, expanding the data size at a lower caption quality does not yield similar gains even at a bigger model ($2^{nd}$ vs $4^{th}$ row), aligning with results reported by Liu et al. (2023c). This highlights that data quality brought by AutoCap is as crucial as the data quantity. Lastly, we examine the effect of scaling the data with synthetic captions. For this, we train for 50k steps by fixing AC and Clotho in the training data and varying the amount of synthetic data. As reported in Fig. 3 (right), increasing the amount of data with synthetic caption has a clear improvement over both IS and FD, with the model trained on the whole AutoReCap achieving the best results.
Model size scaling. In Tab. 5, we report ($1^{st}$ row) that GenAu-L (1.25B params) is constantly favoured over GenAu-S (493M params). This is further confirmed by our automatic evaluation in Tab. 3 ($1^{st}$ vs $3^{rd}$ row), where the larger model shows significant improvements across all metrics. The scaling trend is also evident in Fig. 3, which demonstrates a clear correlation between model size and performance in terms of both IS and FD scores.
Model architecture ablation. Until recently, A UNet (Ronneberger et al., 2015) has been the most popular choice for the diffusion backbone. Yet, as reported in Tab. 4, replacing the FiT backbone with a UNet drastically reduces performance across all metrics. This supports baseline findings where UNet-based methods lag behind transformer-based approaches (Huang et al., 2023a). Another choice that has recently gained popularity is the DiT architecture (Peebles & Xie, 2023). Make-an-Audio-2 (MAD-2) employs a DiT at a similar model size and data scale as GenAU-L. However, as we show in Tab. 5, our model is consistently preferred over MAD-2 ($3^{rd}$ row), even without dataset recaptioning ($4^{th}$ row) (i.e. at similar data settings). We infer that the FiT architect, with its read and write operations, allocates compute more efficiently to the key segments of the input, making it more suitable to ambient audio clips which often include silent or redundant parts.
@@ 5 Conclusion
We take a holistic approach to improve the quality of existing audio generators. Starting by addressing the scarcity of large-scale captioned audio datasets, we build a state-of-the-art audio captioning method, AutoCap, which leverages audio metadata to collect a dataset of 57M annotated audio clips. We then built a latent diffusion model based on a scalable transformer architecture which we trained on our re-captioned dataset to obtain GenAu, a state-of-the-art open-sources model for audio generation. Our approach not only improves ambient audio generation but also opens up possibilities for extending GenAu to other domains, such as speech and music generation. As an additional contribution, we built AutoReCap-XL, a text-audio-video ambient audio dataset with an unprecedented size of $57M$ pairs. AutoReCap-XL can potentially serve as a joint text-audio-video dataset and broadens novel applications such as text-to-audio-video joint generation, a more natural and desired choice of video generation. We leave larger analyses on this dataset for future work.
Appendix
Appendix A AutoReCap-XL Details
This section outlines the collection and filtering processes for AutoReCap-XL.
A.1 Stage 1: Data Selection
We selected existing video datasets primarily from YouTube for the ease of accessing automatic transcriptions. Specifically, we chose 73 million videos from the datasets AudioSet (Gemmeke et al., 2017), VGGSound (Chen et al., 2020), ACAV100M (Lee et al., 2021), VideoCC (Nagrani et al., 2022), YTTEMP1B (Zellers et al., 2022), and HDVila-100M (Xue et al., 2022). We select these datasets for their likelihood of containing videos with strong audio-video correspondence.
A.2 Stage 2: Speech and Music Filtering
We downloaded English transcripts from YouTube and used automatically generated ones for videos without existing transcripts. However, we discard videos without any transcripts. While some datasets provide only video segments with specific timestamps, we processed the full videos, totaling around 73 million videos. We accepted audio segments longer than one second that lacked any corresponding subtitles, indicating the absence of speech and music. After filtering, we isolated approximately 327.3 million segments from 55.1 million videos. Fig. 5 displays the distribution of the number of segments per video. We denote this dataset as AutoReCap-XL-Raw. Subsequently, we use AutoCap to caption the audio segments. Fig. 7 shows the distribution of caption lengths. Given that AutoReCap was trained for 10-second audio, we limited segments to this duration. Additionally, we concatenate consecutive segments yielding identical captions to form longer audio clips. Fig. 9 illustrates the audio length distribution, and a word cloud of the captions is shown in Fig. 11. Despite filtering, the dataset was still dominated by captions related to speech and music. We attribute this to the limitations of YouTube’s automatic transcription, particularly with certain types of music and less common languages.
Figure 4: Distribution of number of segments per-video in AutoReCap-XL-Raw
Figure 5: Distribution of the number of segments per-video in AutoReCap-XL
Figure 6: Distribution of caption length of AutoReCap-XL-Raw
Figure 7: Distribution of caption length of AutoReCap-XL
Figure 8: Distribution of audio duration of AutoReCap-XL-Raw
Figure 9: Distribution of audio duration of AutoReCap-XL
Figure 10: Word cloud of audio captions in AutoReCap-XL-Raw
Figure 11: Word cloud of audio captions in AutoReCap-XL
A.3 Stage 3: Post-filtering of Speech and Music.
To further refine the dataset from speech and music, We follow a simple filtering approach. Specifically, we employed a large language model (LLM) to generate keywords associated with speech and music, such as "talking", "speaking", and "singing," and excluded all audio segments whose captions contained such keywords. This process yielded 57 million audio-text pairs from 20.3 million videos. Fig. 5 shows the number of segments per video, Fig. 7 shows the caption length distribution, Fig. 9 shows the audio length distribution, and Fig. 11 presents a word cloud of the final captions. We outline the data sources for constructing this dataset in Tab. 6. Our proposed dataset is not only 90 times larger than the previously largest available dataset, LAION-Audio-630KWu et al. (2023) in terms of the number of audio clips, but also provides more accurate captions compared to existing datasets that rely on raw textual data. A comprehensive comparison with other datasets is detailed in Tab. 7
Appendix B Limitations.
B.1 AutoCap
Sounds emitted by various objects can often sound similar, such as a waterfall compared to heavy rain, or a can versus a motorcycle engine. In scenarios where metadata lacks detail, our audio captioning model may struggle to disambiguate these sounds accurately. The model also tends to falter in capturing the temporal relationships between sounds and differentiating foreground from background noises. Additionally, since it is fine-tuned on AudioCaps, which contains a limited vocabulary of 4,892 unique words (excluding common stop words), the model frequently produces repetitive words and captions.
B.2 GenAu
@@
Although our model is trained to generate natural sound effects, it underperforms in specialized areas like music generation or text-to-speech synthesis, where more targeted models are superior. Moreover, the limited vocabulary of the paired texts, even though extensive, hampers the model’s ability to accurately generate audio for long and detailed prompts.
B.3 AutoReCap-XL
Our proposed dataset, AutoReCap-XL, is substantial in size but features a constrained vocabulary of only 4,461 unique words, excluding stop words, due to the vocabulary limitations of the AudioCaps-trained captioner. Furthermore, despite its potential as a significant contribution, this dataset has not yet been extensively analyzed for caption accuracy or performance in downstream tasks.
Appendix C Evaluation Details
C.1 Audio Captioning
@@
While the established practice in the evaluation of audio captioning methods is to report the results on the test set using the checkpoint that performs best on the validation subset, prior work (Étienne Labbé et al., 2023; Kim et al., 2024b) reported high instability of the metrics on the validation subset and weak correlation between the validation and test performance, making the model’s results vary significantly for different seeds. To alleviate this, ENCLAP (Kim et al., 2024b) selects around five best-performing validation checkpoints and reports their best results on the test set. CNext-trans (Étienne Labbé et al., 2023) uses the FENSE score to pick the best validation checkpoint. This method of choosing the best checkpoint may produce misleading results and potentially disadvantage baselines. Our model, thanks to the two-stage training paradigm, significantly reduces this instability and we observe steady performance gains as training progresses. Therefore, we report the results at convergence, specifically after 20 epochs of pre-training and 20 epochs of fine-tuning.
C.2 Audio Generation
There is a lack of consistency in the metrics used across text-to-audio generation baselines. Some baselines, such as Liu et al. (2023b) and Huang et al. (2023a), employ the VGGish network (Hershey et al., 2017) to compute the Fréchet Distance, while others, like Liu et al. (2023c), utilize the PANNs network (Kong et al., 2019), and still others rely on OpenL3 embeddings, such as Evans et al. (2024b). Additionally, some baselines use the LAION CLAP network (Wu et al., 2023) to compute the CLAP score, whereas others use the Microsoft CLAP network (Elizalde et al., 2023). To further complicate matters, different baselines often report varying results in various publications. To address these discrepancies, we recalculated all metrics for the baselines using their publicly released checkpoints under identical evaluation configurations. Our method significantly outperforms the baselines across all metrics, except for the Fréchet Distance, where it is slightly behind Make-An-Audio 2 (Huang et al., 2023a). Nevertheless, our user study, detailed in the main paper, indicates that GenAu is generally preferred over Make-An-Audio 2.
Figure 12: A screenshot of the user study interface.
Data Source ##\## pairs
AudioSet 339.4k
VGGSounds 126.9k
Freesounds 262.3k
BBC Sound Effects 31.2k
YouTube Videos 57.0M
   ACAV-100M
   VideoCC
   YTTEMP1B
   HDVila-100M
AutoReCap 761.1k
AutoReCap-XL 57.0M
Table 7: Comparative overview of the main audio-language datasets.
Dataset # Text-Audio Pairs Duration (h) Text source
AudioCaps 52,904 144 Human
Clotho 5,929 37 Human
MACS 3,537 10 Human
WavText5K 4,072 23 Online raw-data
SoundDescs 32,979 1,060 Online raw-data
LAION-Audio-630K 633,526 4,325 Online raw-data
WavCaps 403,050 7,567 Processed raw-data
AutoReCap 761,113 8,763 Automatic re-captioning
AutoReCap-XL 57M 123,500 Automatic re-captioning
AutoReCap-XL-Raw 327.3M - Automatic re-captioning
Table 8: Audio Evaluation Criteria
Criterion Description
Realism Which audio is more realistic? In other words, is more likely to be a result of a real action.
Quality Which audio has better quality, regardless of the realism of the audio. Please note that some audio may have background noise, which should not be confused with low quality.
Prompt Alignment Considering the prompt to generate the audio is "A sewing machine operating as a machine motor hisses loudly in the background", which audio better follows the given prompt?
Overall Preference Considering the realism, quality, and prompt alignment of the audio, which audio do you prefer more overall? The prompt is: "A sewing machine operating as a machine motor hisses loudly in the background."
C.3 User Study
Each user study reported in this paper involved 5 different participants, yielding a total of 1000 responses per study. Samples were selected from the AudioCaps test split, specifically choosing the top 200 samples with the longest text prompts and sampling 50 for each study to enhance the likelihood of obtaining more complex audio scenarios. To minimize discrepancies between baselines, we fix the seed and other sampling parameters across all experiments.
During the user study, participants were initially presented with two audio clips from the compared baselines and asked to judge which one sounded more realistic. They were then prompted to choose the audio they believed had better quality. Next, after showing the prompt used to generate the audio, participants were asked to select the clip that most faithfully followed the prompt. Finally, they were asked to choose their overall preferred audio clip. A screenshot of the user study interface is included in Fig. 12, and the questions posed to the annotators are detailed in Tab. 8.
@@ C.3 User Study
CoNeTTE A man is speaking and wind is blowing
ENCLAP A man is speaking and wind is blowing
@@
CoNeTTE A man speaking with birds chirping in the background.
ENCLAP Birds are chirping and a man speaks
@@
CoNeTTE A telephone ringing followed by a beep.
ENCLAP A telephone dialing followed by a series of electronic beeps
@@
CoNeTTE A train horn blows and a steam whistle is blowing
ENCLAP A train running on railroad tracks followed by a train horn blowing as wind blows into a microphone
@@
CoNeTTE A woman speaking followed by a door opening and closing.
ENCLAP A young girl speaks followed by a loud bang
@@
CoNeTTE A baby is laughing and people are talking.
ENCLAP A baby laughs and splashes in water
@@
CoNeTTE A dog is barking and a person is walking.
ENCLAP Birds chirp and a dog barks
@@
CoNeTTE A toilet is flushed and water is running.
ENCLAP A faucet is turned on and runs
@@
Appendix D Training and Inference Details
D.1 AutoCap
We train the audio captioning model using the Adam optimizer, starting with a learning rate of $10^{-4}$ in stage 1, and reducing to $10^{-5}$ in stage 2. The training was completed over 9 hours on eight A100 80GB GPUs. Although our model is training with 10-second audio clips, we observed qualitatively that it generalizes well to short audios, such as 1-2 second audio clips.
D.2 GenAu
We employ the LAMB optimizer for our audio generation model, setting the learning rate at $0.005$ with a cosine schedule, and incorporating a weight decay of $0.1$ and a dropout rate of $0.1$. The small model variant is trained for 210k steps with a batch size of 2,048, while the large model variant is trained for 220k steps with a batch size of 3,072. The large model is trained over 48 hours on 48 A100 80GB GPUs, and the small model on 32 GPUs. Ablation studies are conducted on eight A100 80GB GPUs using a batch size of 512. We further condition the model on the training dataset with a conditioning dataset ID. For generation, we utilize the AudioCaps dataset ID as it is the most reliable dataset.
Appendix E Additional Results
In this section, we present additional results which are complemented by our Website.
E.1 Additional Audio Captioning Evaluation
In Tab. 9 we show qualitative results of the captions produced by our method and compare them with state-of-the-art AAC methods. See the Website for qualitative results accompanied by the original audio. While ENCLAP (Kim et al., 2024b) and CoNeTTE (Étienne Labbé et al., 2023) tend to produce short captions, our method produces the most descriptive captions, capturing the most amount of elements from the ground truth audio, an important capability to allow high-quality audio generation (Shi et al., 2020).
E.2 Additional Audio Generation Evaluation
In Tab. 11, we ablate the 1D-VAE bottleneck size in terms of reconstruction loss and performance of a subsequently trained latent audio diffusion model, in terms of FAD, FD, and IS. Similarly to the phenomenon observed in the image and video generation domain (Gupta et al., 2023; Esser et al., 2024), we observe that a larger number of channels allocated to the latent space results in lower reconstruction losses, but making the latent space more complex, hindering generation quality. We adopt 64 1D-VAE channels for all our experiments.